uzsbgck:

5.93 cubic centimetres to cubic metres ( Use Scientific notation)

AZ:

1 m = 100 cm 1m^3 = 1000000 cm^3

AZ:

So first, can you convert 5.93 cubic centimeters to cubic meters?

xxemilianaxx:

i dont think there gonna answer

uzsbgck:

How do I do that @AZ

AZ:

Just divide by 1,000,000 5.93 cm^3 * (1 m^3 /1,000,000 cm^3) = ?? m^3

uzsbgck:

so 0.00000593

AZ:

Yes so we have 

AZ:

How many times do we move it to the right?

AZ:

That's going to be the number on our exponent for 5.93 * 10^(that number)

AZ:

Also, it's going to be a negative number because we're moving it to the right

uzsbgck:

6 times?

AZ:

yes if we were moving 6 times the other way then it would just be 10^6 but since we're moving 6 times towards the right, it's 10^(-6) And to make sense of it- just think about it. When we have 10^6. We know it's going to be a really big number 10^(-6) means we have a really small number

uzsbgck:

ok so would the whole number be 5.93 because thats where it was moved 2

AZ:

Also, we didn't exactly have to do all this math. Remember what we started out with? \( 5.93 \text{cm}^3 \times \dfrac{1 \text{m}^3}{1,000,000 \text{cm}^3} = ??~ \text{m}^3\) what is another way to write 1,000,000? It's 10^6! and \(10^6\) is the same thing as \( 10^{-6}\) because \(\dfrac{1}{a^b} = a^{-b}\)

AZ:

Yes!

AZ:

So your final answer is \(5.93 \times 10^{-6}\)